# Multichannel Conflict-Avoiding Codes for Expanded Scenarios

### v1

@@ VII-A At Most One-Packet Per Time Slot MC-CACs
The following upper bound of $A(M,L,w)$ is due to [4, Corollaries 3.6, 3.7].
@@
The upper bounds in Theorem 16 can be improved in the case where $M<w$, which is shown in the following theorem.
@@
We remark that an alternate proof of Theorem 16 can be obtained by means of the arguments in Appendix A.
@@ VII-B Mixed-Weight Codes
[5] introduced mixed-weight CACs for the purpose of increasing the throughput and reducing the access delay of some potential users with higher priority. This section will provide some constructions of mixed-weight CACs, which can be seen as generalizations of Theorem 6 and Theorem 9. Furthermore, the concept of mixed-weight will also be extended to a multichannel version by relaxing the identical-weight constraint in prior studies of MC-CACs.
@@
We can use the method given in [5, Theorem 16] to construct a mixed-weight CAC of length $(w-1)p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}$ with weight-set $\mathcal{W}=\{w-1,w,w_{i_{1}}^{*},w_{i_{2}}^{*},\ldots,w_{i_{t}}^{*}\}$ provided a code in $\mathrm{CAC^{e}}(p_{i_{j}}^{r_{i_{j}}},w_{i_{j}}^{*})$ exists for some $i_{1},i_{2},\ldots,i_{k}$.
@@
Let $w,r_{1},r_{2},\ldots,r_{n}$ be positive integers and $p_{1}<p_{2}<\cdots<p_{n}$ be primes such that $p_{1}\geq w$. Suppose, for each $1\leq i\leq n$, the two conditions given in (26) and (27) hold and there exists a code $\mathcal{A}_{i}\in\mathrm{CAC}(p_{i}^{r_{i}},w_{i}^{*})$ that contains $m_{i}$ equi-difference codewords, where $w_{i}^{*}$ is an arbitrary positive integer and $m_{i}$ is not necessarily non-zero. Then there exists a code $\mathcal{C}\in\mathrm{CAC}(L,\mathcal{W})$ with
$|\mathcal{C}|=\frac{p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}-1}{2}+1+\sum_{i=1}^{n}m_{i},$
where $L=(w-1)p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}$ and $\mathcal{W}=\{w-1,w,w_{1}^{*},w_{2}^{*},\ldots,w_{n}^{*}\}$. In particular, if all codewords in each $\mathcal{A}_{i}$ are non-exceptional, then $\mathcal{C}$ contains $1+\sum_{i=1}^{n}m_{i}(w_{i}^{*}-1)$ codewords with weight w1w-1, $\frac{p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}}{2}-\sum_{i=1}^{n}m_{i}(w_{i}^{*}-1)$ codewords with weight $w$, and $m_{i}$ codewords with weight $w_{i}^{*}$ for each $i$.
@@
Let $p_{1}<\cdots<p_{n}$ be primes and $w_{1},\ldots,w_{n}$ be positive integers (not necessarily distinct) such that $p_{i}\geq 2w_{i}-1$ for any $i$. If there is a code in $\mathrm{CAC^{e}}(p_{i},w_{i})$ with $m_{i}$ codewords for each $i$, then for any positive integers $r_{1},\ldots,r_{n}$, there exists a mixed-weight code $\mathcal{C}\in\mathrm{CAC}(L,\mathcal{W})$ with
$|\mathcal{C}|=\sum_{i=1}^{n}\frac{m_{i}(p_{i}^{r_{i}}-1)}{p_{i}-1}\prod_{j=i+1}^{n}p_{j}^{r_{j}},$
where $L=p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}$ and $\mathcal{W}=\{w_{1},\ldots,w_{n}\}$. In particular, $\mathcal{C}$ contains
$\frac{m_{i}(p_{i}^{r_{i}}-1)}{p_{i}-1}\prod_{j=i+1}^{n}p_{j}^{r_{j}}$
codewords with weight $w_{i}$, for $i=1,2,\ldots,n$.
@@
The proof can be done with the same arguments in the proof of Theorem 9 by adjusting $\Gamma_{i}$ being the set of $m_{i}$ generators of the given code in $\mathrm{CAC^{e}}(p_{i},w_{i})$ and replacing the definition of $S_{a}$ in (37) with
We now extend the concept of mixed-weight to multichannel CACs.
@@ VIII Concluding Remarks
It has been shown in [1, Theorem 1] that
@@
under some specific parameter settings. For general $M\geq 3$, when $w>M$ and under other specific parameter settings, we can see from Theorem 14 that the lower bound of $K(M,L,w)/L$ approaches

### v2

@@ VII-A At Most One-Packet Per Time Slot MC-CACs
The following upper bound of $A(M,L,w)$ is due to [30, Corollaries 3.6, 3.7].
@@
Note that Theorem 16 can be derived by the same argument we used in the proof of Theorem 3, which is provided in Appendix A.
The upper bounds in Theorem 16 can be further improved in the case where $M<w$, which is shown in the following theorem.
@@ VII-B Mixed-Weight Codes
Mixed-weight CACs were introduced [20] for the purpose of increasing the throughput and reducing the access delay of some potential users with higher priority. This subsection will provide some constructions of mixed-weight CACs, which can be seen as generalizations of Theorems 6 and 9. Furthermore, the concept of mixed-weight will also be extended to a multichannel version by relaxing the identical-weight constraint in prior studies of MC-CACs.
@@
Let $\mathrm{CAC}(L,\mathcal{W})$ denote the class of all mixed-weight CACs of length $L$ with weight-set $\mathcal{W}$.
Under the assumption that a code in $\mathrm{CAC}(p_{i}^{r_{i}},w_{i}^{*})$ exists for each $i=1,2,\ldots,n$, we will first construct a mixed-weight CAC of length $p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}$ with weight-set and then construct a mixed weight CAC of length $(w-1)p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}$ with weight-set $\mathcal{W}=\{w-1,w_{1}^{*},w_{2}^{*},\ldots,w_{n}^{*}\}$. Note that the two constructions are obtained by modifying those in Theorems 9 and 6, respectively.
@@
Let $w_{i},r_{i}$, $1\leq i\leq n$, be positive integers. Assume that $p_{1},p_{2},\ldots,p_{n}$ are primes such that $p_{i}\geq 2w_{i}-1$ for each $i$. Suppose, for each $1\leq i\leq n$, there exists a code $\mathcal{A}_{i}\in\mathrm{CAC}(p_{i}^{r_{i}},w_{i})$ that contains $m_{i}$ equi-difference codewords. Then there exists a code $\mathcal{C}\in\mathrm{CAC}(L,\mathcal{W})$ with
$|\mathcal{C}|=\sum_{i=1}^{n}m_{i}\prod_{j=i+1}^{n}p_{j}^{r_{j}},$
where $L=p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}$ and $\mathcal{W}=\{w_{i}\}_{i=1}^{n}$, that contains $m_{i}\prod_{j=i+1}^{n}p_{j}^{r_{j}}$ codewords of weight $w_{i}$ for each $i$.
Proof.
Let $\Gamma_{i}$ be the set of $m_{i}$ generators of the equi-difference codewords of $\mathcal{A}_{i}$. The proof can be done with the same arguments as in the proof of Theorem 9 by defining
$\widehat{\Gamma}_{i}\triangleq\{(0,\ldots,0,g,x_{i+1},\ldots,x_{n})$ $\in\mathbb{Z}_{p_{1}^{r_{1}}}\times\cdots\times\mathbb{Z}_{p_{n}^{r_{n}}}:\,$
$g\in\Gamma_{i},x_{j}\in\mathbb{Z}_{p_{j}^{r_{j}}}\text{ for }i<j\leq n\}$
and replacing the definition of $S_{a}$, for $a\in\widehat{\Gamma}_{i}$, in (40) with
The rest of the proof is omitted. ∎
@@
Let $w_{i},r_{i}$, $1\leq i\leq n$, and $w$ be positive integers. Assume that $p_{1},p_{2},\ldots,p_{n}$ are primes such that $p_{i}\geq 2w-1,2w_{i}-1$ for each $i$. Suppose, for each $1\leq i\leq n$, the two conditions given in (27) and (28) hold and there exists a code $\mathcal{A}_{i}\in\mathrm{CAC}(p_{i}^{r_{i}},w_{i})$ that contains $m_{i}$ equi-difference codewords. Then there exists a code $\mathcal{C}\in\mathrm{CAC}(L,\mathcal{W})$ with
$|\mathcal{C}|=\frac{p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}+1}{2}+\sum_{i=1}^{n}m_{i}\prod_{j=i+1}^{n}p_{j}^{r_{j}},$
where $L=(w-1)p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}$ and $\mathcal{W}=\{w-1\}\cup\{w_{i}\}_{i=1}^{n}$. In particular, $\mathcal{C}$ contains $(p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}+1)/2$ codewords of weight w1w-1 and $m_{i}\prod_{j=i+1}^{n}p_{j}^{r_{j}}$ codewords of weight $w_{i}$ for each $i$.
@@
We adopt the notation and definitions used in the proof of Theorem 6. Let $L^{\prime}=p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}$ for convenience. Note that $\mathbb{Z}_{L}\cong\mathbb{Z}_{w-1}\times\mathbb{Z}_{L^{\prime}}$ and $\mathbb{Z}_{L^{\prime}}\cong\mathbb{Z}_{p_{1}^{r_{1}}}\times\cdots\times\mathbb{Z}_{p_{n}^{r_{n}}}$. We aim to construct $n+1$ classes of codewords, say $\mathcal{C}_{w-1},\mathcal{C}_{w_{1}},\ldots,\mathcal{C}_{w_{n}}$, consisting of codewords with weights $w-1,w_{1},\ldots,w_{n}$, respectively.
Firstly, for $a\in\widehat{Q}$, define
$S_{a}=\{j(1,a)\in\mathbb{Z}_{w-1}\times\mathbb{Z}_{L^{\prime}}:\,j=0,1,2\ldots,w-2\}$
be the (w1)(w-1)-subset generated by $a$, where $\widehat{Q}$ is defined in (30). It follows from the proof of Theorem 6 that $d^{*}(S_{a})$, $a\in\widehat{Q}$, are mutually disjoint. Note that only Case 2 in the proof of Theorem 6 is needed here. Moreover, we have
$\bigcup_{a\in\widehat{Q}}d^{*}(S_{a})\subseteq\mathbb{Z}_{w-1}^{*}\times\mathbb{Z}_{L^{\prime}}^{*}.$ (78)
Then, define the codeword $T=\{(j,0)\in\mathbb{Z}_{w-1}\times\mathbb{Z}_{L^{\prime}}:\,j=0,1,\cdots w-2\}$. Obviously, $d^{*}(T)=\mathbb{Z}_{w-1}^{*}\times\{0\}$, and it is disjoint from $d^{*}(S_{a})$ for every $a\in\widehat{Q}$. Let $\mathcal{C}_{w-1}=\{T\}\cup\{S_{a}:\,a\in\widehat{Q}\}$. It follows from (78) that
$\bigcup_{S\in\mathcal{C}_{w-1}}d^{*}(S)\subseteq\mathbb{Z}_{w-1}^{*}\times\mathbb{Z}_{L^{\prime}}.$ (79)
By (30), we therefore have $|\mathcal{C}_{w-1}|=1+(L^{\prime}-1)/2$.
Finally, let $\mathcal{C}_{w_{i}}$ be the sets of codewords of weight $w_{i}$ that are obtained from $\mathcal{A}_{i}$ by the construction given in Theorem 20 and raised from subsets of $\mathbb{Z}_{L^{\prime}}$ to $\{0\}\times\mathbb{Z}_{L^{\prime}}$. More precisely, each codeword in $\mathcal{C}_{w_{i}}$ is of the form $\{(0,s)\in\mathbb{Z}_{w-1}\times\mathbb{Z}_{L^{\prime}}:\,s\in S\}$, where $S$ is a codeword obtained from $\mathcal{A}_{i}$ by the construction given in Theorem 20. It is clear that
$\bigcup_{S\in\mathcal{C}_{w_{1}}\cup\cdots\cup\mathcal{C}_{w_{n}}}d^{*}(S)\subseteq\{0\}\times\mathbb{Z}_{L^{\prime}}.$
Hence, it follows from (79) and Theorem 20 that $\mathcal{C}_{w-1}\cup\mathcal{C}_{w_{1}}\cup\cdots\cup\mathcal{C}_{w_{n}}$ forms the desired mixed-weight code. ∎
Example 5.
Let $w=4$, $n=2$, $r_{1}=r_{2}=1$ and $p_{1}=23,p_{2}=47$. We have $L=3243$. Consider $w_{1}=9,w_{2}=6$ and $m_{1}=1,m_{2}=2$, and assume that the set of generators of equi-difference codewords in $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ are $\Gamma_{1}=\{1\}$ and $\Gamma_{2}=\{1,7\}$, respectively.
By Theorems 6 and 21, we have $\widehat{Q}=\widehat{Q}_{1}\cup\widehat{Q}_{2}$, where $\widehat{Q}_{1}=\{(g,x)\in\mathbb{Z}_{23}\times\mathbb{Z}_{47}:\,g\in Q_{1},x\in\mathbb{Z}_{47}\}$, $\widehat{Q}_{2}=\{(0,g)\in\mathbb{Z}_{23}\times\mathbb{Z}_{47}:\,g\in Q_{2}\}$, $Q_{1}=\{1,2,3,4,6,8,9,12,13,16,18\}$, and $Q_{2}=\{1,2,3,4,6,7,8,9,12,14,16,17,18,21,24,25,27,28,32,34,36,37,42\}$. Let $\mathcal{C}_{3}=\{T\}\cup\{S_{a}:\,a\in\widehat{Q}\}$, where $T=\{(0,0),(1,0),(2,0)\}$ and $S_{a}=\{(0,0),(1,a),(2,2a)\}$. Note that the elements in those codewords are considered in $\mathbb{Z}_{3}\times\mathbb{Z}_{1081}$, and can be mapped to elements in $\mathbb{Z}_{3243}$ by the CRT correspondence (4). For example, $T=\{0,1081,2162\}$ and $S_{(1,0)}=\{0,2209,1175\}$.
Following the construction in Theorems 20 and 21, we have $\mathcal{C}_{9}=\{S_{(1,x)}:\,x=0,1,\ldots,46\}$ and $\mathcal{C}_{6}=\{S_{(0,1)},S_{(0,7)}\}$, where
$S_{(1,x)}=\{(0,j,jx)\in\mathbb{Z}_{3}\times\mathbb{Z}_{23}\times\mathbb{Z}_{47}:\,j=0,1,\ldots,8\}$
and
$S_{(0,k)}=\{(0,0,jk)\in\mathbb{Z}_{3}\times\mathbb{Z}_{23}\times\mathbb{Z}_{47}:\,j=0,1,\ldots,5\}$
for $k=1,7$. Similarly, by the CRT correspondence $\mathbb{Z}_{3}\times\mathbb{Z}_{23}\times\mathbb{Z}_{47}\cong\mathbb{Z}_{3243}$, we have $S_{(0,1)}=\allowbreak\{0,897,1035,1932,2070,310\}$, $S_{(0,7)}=\{0,552,759,1518,2277,3036\}$, and $S_{(1,0)}=\{0,141,\allowbreak 282,1128,1269,1410,2256,2397,2538\}$ for instance. Hence we have $|\mathcal{C}_{3}|=541$ , $|\mathcal{C}_{9}|=47$, $|\mathcal{C}_{6}|=2$.
We now extend the concept of mixed-weight to MC-CACs.
@@
Let $\text{MC-CAC}(M,L,\mathcal{W})$ denote the class of all mixed-weight MC-CACs of length $L$ with weight-set $\mathcal{W}$ for $M$ channels.
The construction of MC-CACs in Theorem 11 relies on the CAC constructions in Theorems 6 and 9. Since these admit mixed-weight generalizations (Theorems 21 and 20, respectively), Theorem 11 extends naturally to the mixed-weight case.
Theorem 22.
Let $w_{i},r_{i}$, $1\leq i\leq n$, and $w$ be positive integers. Assume that $p_{1},p_{2},\ldots,p_{n}$ are primes such that $p_{i}\geq 2w-1,2w_{i}-1$ for each $i$. Suppose, for each $1\leq i\leq n$, the two conditions given in (27) and (28) hold, and there exists a code in $\mathrm{CAC}(p_{i}^{r_{i}},w_{i})$ that contains $m_{i}$ equi-difference codewords. Then there exists a mixed-weight code $\mathcal{C}\in\mathrm{CAC}(2,L,\mathcal{W})$ with
$|\mathcal{C}|=1+\prod_{i=1}^{n}p_{i}^{r_{i}}+\sum_{i=1}^{n}\frac{2m_{i}(p_{i}^{r_{i}}-1)}{p_{i}-1}\prod_{j=i+1}^{n}p_{j}^{r_{j}}$
where $L=(w-1)p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}$ and $\mathcal{W}=\{w,w-1\}\cup\{w_{i}\}_{i=1}^{n}$. In particular, $\mathcal{C}$ contains $\prod_{i=1}^{n}p_{i}^{r_{i}}$ codewords of weight $w$, one codeword of weight w1w-1, and $\big(2m_{i}(p_{i}^{r_{i}}-1)/(p_{i}-1)\big)\prod_{j=i+1}^{n}p_{j}^{r_{j}}$ codewords of weight $w_{i}$, for $i=1,2,\ldots,n$.
Proof.
The proof follows from the combination of the proofs of Theorems 11 and 20. The codeword of weight w1w-1 is given by $\{(2,(j,0)):\,j=0,1,\ldots,w-2\}$, which is a variation of $U$, the last codeword in the proof of Theorems 11. ∎
We also have the following mixed-weight version of Theorem 14.
Theorem 23.
Let $M,L,L^{\prime}$, and $w$ be positive integers such that $w>M\geq 3$, $M$ divides $w$, $L=(2\frac{w}{M}-1)L^{\prime}$, and all prime factors of $L^{\prime}$ are larger than or equal to 2w12w-1. If, for $1\leq m\leq M$, there is a code $\mathcal{C}_{m}\in\mathrm{CAC}(L^{\prime},\mathcal{W}_{m})$ with $|\mathcal{C}_{m}|=h_{m}$ codewords, then there exists a code $\mathcal{C}\in\text{MC-CAC}(M,L,\mathcal{W})$ with $|\mathcal{C}|=L^{\prime}+\sum_{m=1}^{M}h_{m}$ codewords, where $\mathcal{W}=\{w\}\cup\bigcup_{m=1}^{M}\mathcal{W}_{m}$ and each $\mathcal{W}$ is an arbitrary set of positive integers. In particular, $\mathcal{C}$ contains $L^{\prime}$ codewords of weight $w$, and the number of codewords of weight $w^{\prime}$ in $\mathcal{C}$ equals that in $\bigcup_{m=1}^{M}\mathcal{C}_{m}$.
Proof.
The proof can be obtained by a straightforward adaptation of the proof of Theorem 14. Specifically, the codewords on channel $m$ are constructed using $\mathcal{C}_{m}$. That is, for $S\in\mathcal{C}_{m}$, $1\leq m\leq M$, let $\widehat{S}=\{(m,(0,a))\in\mathcal{I}_{M}\times\mathbb{Z}_{L}:\,a\in S\}$. ∎
@@ VIII Concluding Remarks
This paper investigates MC-CACs from the perspectives of both constructions and theoretical upper bounds on the maximum code size. We first introduce the notion of exceptional codewords for MC-CACs and establish a series of structural properties using techniques from additive combinatorics, including Kneser’s theorem. These results yield new upper bounds on the maximum code size of MC-CACs, especially for the case where the channel number is less than the maximum number of active users. Next, we generalize two previously known constructions of (single-channel) CACs in Theorems 6 and 9, which in turn lead to a construction of MC-CACs with $M=2$ channels, as presented in Theorem 11. Necessary conditions for these constructions to be optimal are characterized in Theorems 7, 10, and 13, respectively. A construction of MC-CACs for $M\geq 3$ such that $M$ divides $w$ is given in Theorem 14, and a corresponding upper bound on the maximum code size is established in Theorem 15. Our results on MC-CACs can be further extended to AM-OPPTS MC-CACs in Theorems 1719. Finally, we study mixed-weight MC-CACs for the first time and propose a series of constructions in Theorems 22 and 23, based on the structures of mixed-weight CACs given in Theorems 21 and 20.
It has been shown in [28, Theorem 1] that
@@
under some specific parameter settings. For general $M\geq 3$, when $w>M$ and under other specific parameter settings, we can see from Corollary 2 that the lower bound of $K(M,L,w)/L$ approaches